OpenStudy (anonymous):

The general formula for kinetic energy is k = 0.5mv^2 where k = kinetic energy, m = mass, and c = velocity. Solve k = 0.5mv^2 for v.

OpenStudy (anonymous):

any variables given?

OpenStudy (anonymous):

Naw there uh answer choices though I forgot to put that on here v = m/2k V = square root of the quantity m/2k V = 2k/m V = square root of the quantity 2k/m

OpenStudy (anonymous):

\[k=0.5mvv^2\] if you solve with simple algebra you should be able to get the answer